A rectangular billboard 5 feet in height stands in a field so that its bottom is 6 feet above the ground. A nearsighted cow with eye level at 4 feet above the ground stands x

x

 feet from the billboard. Express θ

θ

, the vertical angle subtended by the billboard at her eye, in terms of x

x

. Then find the distance x


x


 the cow must stand from the billboard to maximize θ

θ

.


1
Expert's answer
2022-01-10T16:06:03-0500


From the diagram above, we can have that

"CD=x ," which is the distance of the base.

"B_1D=1, B_2D=7,"

"\theta=\beta-\alpha"

"\theta=\tan^{-1}(\frac{7}{x})-\tan^{-1}(\frac{1}{x})"

Find the derivative of "\theta" wrt "x"

"\frac{d \theta}{dx}=-\frac{7}{x^2+49}+\frac{1}{x^2+1}"

Set the derivative to 0

"-\frac{7}{x^2+49}+\frac{1}{x^2+1}=0\\
\text{Multiply through by }(x^2+49)(x^2+1)\\
-7(x^2+1)+(x^2+49)=0\\
-7x^2-7+x^2+49=0\\
-6x^2+42=0\\
x^2=7\\
x=\sqrt{7}"

Thus the critical point is "\sqrt{7}" . Let check if "x=\sqrt{7}" is a maximum point

"\theta'(\sqrt{7})=-\frac{3}{56}\sqrt{7}<0" .

Thus, "x=\sqrt{7}" is a maximum point.

Hence the distance of the cow from the billboard must be "\sqrt{7}ft" so as to maximize it